OpenStudy (brian.sama):

If f(x) = 8 – 10x and g(x) = 5x + 4, what is the value of (fg)(–2)?

TheSmartOne (thesmartone):

Is that \( (f \cdot g)(x)\) or \((f ~o~g)(x)\)?

TheSmartOne (thesmartone):

I just want to make sure, since most kids around here copy paste their questions, and it gets messed up. :P If it is \( (f \cdot g)(x)\) then \( (f \cdot g)(x) = f(x) \times g(x)\) If however, it is\( (f~o~g)(x)\) then \( (f~o~g)(x)= f[g(x)]\) You basically substitute g(x) in place of the x in f(x)